OpenStudy (anonymous):

derivative of cos(1/x) / (1 + cosx) (just wanna check my work)

OpenStudy (anonymous):

((1+cosx)(-sin(1/x)(-1x^-2)-cos(1/x)-sin(x))/(1+cosx)^2

OpenStudy (anonymous):

\[\frac{(1 + cosx)(-\sin(\frac{1}{x})(-\frac{1}{x^2}) - (\cos(\frac{1}{x})(-sinx)}{(1 + cosx)^2}\]

OpenStudy (anonymous):

Then what ever that is simplified. If you want to check your work i suggest a site called www.wolframalpha.com.

OpenStudy (anonymous):

wolfram is AMAZING! mathamatica is so good!

OpenStudy (anonymous):

Scroll down and you will see the derivative.